OpenStudy (anonymous):

can someone help me? Given f(x) = 4x^2+3x-5 and g(x)= -2x+12, Find (fg)(x)

OpenStudy (anonymous):

You plug in g(x) into f(x). \[(f\circ g)(x)\]\[= f(g(x)) \]\[= 4(g(x))^2 + 3(g(x)) - 5\]\[=4(-2x+12)^2 + 3(-2x+12) - 5\]\[=4(4x^2 - 48x + 144) + 3(-2x + 12) - 5\]\[=\ ...\]

OpenStudy (anonymous):

Ok. Thank you. I'll try to work the problem now.